A car moves from rest accelarating at 4m/s within 4 sec during the next 10 sec the car moves uniformly after what they break and the car slow down at 8m/s until it stops. find the total distance moved by the car


1
Expert's answer
2020-01-23T06:09:22-0500

1.


"d_1=0.5a_1t_1^2=0.5(4)(4)^2=32\ m"

"v=a_1t_1=(4)(4)=16\frac{m}{s}"

2.


"d_2=vt_2=(16)(10)=160\ m"

3.


"v^2-0^2=2a_2d_3"

"16^2=2(8)d_3"

"d_3=16\ m"

So, the total distance moved by the car:


"d=32+160+16=208\ m"